# How to Build a Quantum Supercomputer: Scaling Challenges and Opportunities

### v1

@@ I.2.3 Challenges at 10k–100k physical qubits
At the very large scale, circuit-level scaling challenges become significant (Markov, 2014, Table 1), including verification, testing, and debugging. For conventional integrated circuits, the challenge of “dark silicon” arises, where a significant fraction of the chip performs various service roles. In quantum computing, FTQC creates a similar overhead.
@@
At the scale of tens of thousands of physical qubits, many fault-tolerant protocols including full-fledged magic state distillation units can be implemented and validated. Yet, the high space and time overhead of FTQC mean this scale will still fall short of demonstrating quantum utility. This prompts the need for advancements in QEC and FTQC schemes that reduce the overhead of fault tolerance, a goal actively pursued in current research trends for building “good” QECCs, i.e., those with high encoding rates, such as the quantum LDPC codes Breuckmann and Eberhardt (2021).
@@
Moreover, the successful realization of QEC requires low-latency integration of QPUs with GPUs, which are very effective in executing a large number of identical, relatively shallow computations in parallel on different input data. Such advantages are relevant to real-time decoding, as discussed in Section III.6. Another approach to relaxing the requirements of decoders is construction of better codes with faster decoding algorithms. It has been speculated that there may exist QECCs whose decoding time is independent of the code distance Gottesman (2022).
Verification, testing, and debugging. As QEC circuits become more sophisticated and undergo optimization to reduce overhead, the possibilities for introducing design errors during these optimizations increase. Verification aims to catch design bugs as soon as they are introduced, testing looks for problematic behaviors in a physical quantum computer (by running specific circuits), and debugging attempts to diagnose and correct problems. Historically, each of these steps became a bottleneck to scaling of classical semiconductor circuits, and required the development of new algorithmic technologies and hardware solutions to sustain scaling Lavagno et al. (2016).
@@
These tasks are much more complicated for quantum circuits than classical ones. For example, a quantum counterpart to the conventional equivalence-checking technique Yamashita and Markov (2010) must tackle unitary operators acting on exponentially large Hilbert spaces. Similarly, testing must be heavily optimized to handle the large number of trials required in view of the non-deterministic nature of quantum measurements and the frequent need of QPUs for recalibration Maksymov et al. (2020). More-sophisticated approximate testing Romanik (1997) techniques must also be developed to take the error tolerance of quantum computation into account. Finally, it is much more complicated to diagnose and eliminate errors Metwalli and Meter (2023); therefore, more-scalable debugging techniques are needed which can benefit from HPC hardware support.
@@ I.2.4 Challenges at 100k–1M physical qubits and beyond
Computation at extreme scales faces system-level and complexity-theoretical challenges as well as those related to physical embedding and distributed computation as quantum interconnects become a bottleneck (Markov, 2014, Table 1). For quantum computers dominated by QEC, these challenges take on specific forms, as described below. Additionally, finding “killer apps” for quantum computers and validating their performance remains challenging.
@@
Distributed FTQC. To achieve utility scale, tens to thousands of logical qubits are needed (Table 3). Even at a 10-4 two-qubit error rate, this translates to 1 million (or more) physical qubits, which is about an order of magnitude more than what can practically be placed in today’s dilution refrigerators (DR). Significantly larger DRs will be costly. Therefore, performing large-scale FTQC will likely require quantum interconnects between multiple DRs. Optical interconnects have been proposed for providing such quantum links between distinct DRs Lee et al. (2020a); Youssefi et al. (2021); Lecocq et al. (2021); Awschalom et al. (2021); Delaney et al. (2022). We discuss the first analysis of the effects of noisy optical interconnects for QECCs on superconducting computers in Section III.7 and discuss the compilation of FTQC algorithms on a multi-DR architecture in Section III.7.
Given the length of utility-scale FTQC algorithms and the need for frequent recalibration of QPUs, a quantum operating system (QOS) must manage an excess supply of on-boarded and off-boarded DRs mid-runtime and dispatch quantum characterization, verification, and validation (QCVV) protocols on them with an appropriate cadence. This reveals a fundamental difference between FTQC compilers and the conventional compilers for classical computers: classical compilers do not require knowledge of the noise profile of the hardware, whereas the choice of QEC codes, their sizes, and ancillary modules (e.g., magic state factories) all require detailed information about the hardware noise characteristics. We describe some preliminary steps toward addressing these challenges for FTQC compilers in Section III.1.
@@
Although it has been hypothesized that one can use optical interconnects between DRs to entangle qubits, the technology is still in its infancy. Therefore, in Section VI we show how adaptive circuit knitting strategies could delay the need for optical interconnects via high-performance distributed quantum evaluation of sub-circuits and classical post-processing to merge the solutions. With large-scale integration of a quantum accelerator with a classical supercomputer, dynamically dispatching workloads to and from the quantum computer will become a complex challenge. We will discuss the use of near real-time circuit synthesis and dispatching in Section V.3 on high-performance quantum–classical workload scheduling.
@@
Micro-architecture standardization. Efficient compilation and execution of large FTQC programs demands optimized and modular instruction-set architectures that are drastically different from that of conventional computers. A challenge at this scale is converging to optimal and standard micro-architectures for the computer. Recent studies Litinski (2019); Beverland et al. (2022); Silva et al. (2024a, b) suggest that efficient instruction pipelines for FTQC using solid-state qubits comprise (potentially multi-level) magic state factories for producing high-quality resource states for consumption in a memory unit (which we call the core processor; see Section III.1). Therefore, unlike the conventional von Neumann architecture wherein data is taken from memory blocks to the gates, in the FTQC pipelines, high-quality gates are brought to the memory block. Thus, FTQC micro-architectures may better resemble that of in-memory computing technologies Verma et al. (2019). This suggests a rethinking of conventional memory hierarchies (e.g., cache, L1, and L2) for quantum memory blocks. Notably, quantum random access memory (QRAM) Giovannetti et al. (2008) is used in quantum computing literature as means for accessing and manipulating classical or quantum data in superposition. However, monumental resources are required to insure its fault tolerance Matteo et al. (2020). Quantum LDPC codes may provide a path forward for realization of feasible quantum memory blocks.
@@
Quantum algorithm discovery. The absence of efficient quantum memory blocks and methods for reading from and writing into them is one of the reasons for the lack of useful quantum algorithms for conventional enterprise problems involving classical big data, for which classical AI has provided major breakthroughs at an increasing pace. This read–write bottleneck eliminates the exponential speedups promised by many algorithms, e.g., quantum machine learning (QML) applied to classical data. Such algorithms include quantum linear system solvers Harrow et al. (2009); Childs et al. (2017), quantum clustering Horn and Gottlieb (2001), quantum principal component analysis Lloyd et al. (2014), and quantum support vector machines Rebentrost et al. (2014). However, the promise for exponential quantum advantage still holds for quantum data Huang et al. (2022) and substantial progress has been made towards overcoming the known trainability limits of quantum neural networks Larocca et al. (2024).
@@
Moreover, processing classical data, e.g., via coherent arithmetic operations, is very costly for quantum computers Häner et al. (2018a, b). Indeed, the qubitized electronic-structure quantum simulation for which we provide resource estimates in Section IV uses look-up tables (coined as quantum read-only memory, or QROM) to avoid calculating trigonometric functions Babbush et al. (2018). In general, applications with classical inputs and outputs are fundamentally limited because quantum computers do not provide a universal advantage. For example, there is provably no quantum advantage for standard comparison-based sorting — no quantum algorithm can solve the task asymptotycally faster than conventional algorithms do Høyer et al. (2001).
Figure 1: Architecture diagram of a quantum–classical full-stack solution. Extensions within the HPC programming environment include a quantum interface library for seamless invocation of quantum kernels, an adaptive circuit knitting hypervisor for efficient quantum workload partitioning and distribution, and quantum compiler and runtime extension for performant quantum circuit compilation. A customized hybrid workload manager ensures maximal quantum resource utilization in a multi-user environment. For fault-tolerant quantum computation, a compiler, emulator, assembler, and real-time decoder work together to use hardware noise profiles to synthesize optimized fault-tolerant circuits. Calibration and control of quantum resources use specialized hardware and are integrated with HPC. At the hardware layer, heterogeneous coprocessors include CPUs/GPUs, quantum processing units (QPUs), probabilistic processing units (PPUs), FPGA and custom-design ASIC with high speed low-latency scale-up interconnect.
@@
Even if a given subroutine in a important application is accelerated by a quantum oracle, it must be a single distinctive bottleneck, otherwise the impact of quantum speedup will be capped by Amdahl’s law Rodgers (1985). Shor’s algorithm for integer factorization and its variants offer a rare combination of likely exponential quantum speedup with a practical need for running the algorithm on many different inputs. In general, quantum optimization could offer a quadratic speedup Szegedy (2004) with many implicit assumptions such as lack of explicit structures in the problem instances. However, the opportunity for such quadratic speedups in practice might be slim and most likely eventually washed away at scale by the huge overhead of QEC Babbush et al. (2021). Engineering at scale requires developing novel quantum heuristic algorithms that work in concert with their classical counterparts, see Section VII for a discussion on the possibility of accelerating classical probabilistic sampling by quantum fluctuations. Recently, there has been tremendous interest on the family of quantum approximate optimization algorithms (QAOA) Farhi et al. (2014); Blekos et al. (2024), including numerical or theoretical studies that claim potential scaling advantages might be possible Shaydulin et al. (2024); Boulebnane and Montanaro (2024); Augustino et al. (2024); Montanaro and Zhou (2024). However, one has to be careful with small-sized effects or the contrived nature of many benchmarking problems that could hinder a true scaling advantage in practical scenarios for which highly tuned classical heuristics are available. Historically, it has proven difficult to develop quantum heuristics that could stay relevant at scale. The search for new quantum algorithms on classical inputs remains a major avenue of research.
@@
Validation of quantum algorithms. Even for quantum mechanical problems for which quantum computation is envisaged to provide revolutionary quantum advantage Lloyd et al. (2014); Huang et al. (2022), the validation of such algorithms—ensuring that they produce correct outputs, especially in practice—remains a challenge. For example, quantum computation of electronic spectra of molecules relies on preparation of input states with significant overlap with the ground state (whose energy is to be estimated). However, is it unclear whether the commonly adopted choices (e.g., the Hartree–Fock state) will be sufficient. Indeed, ground state preparation is a QMA-hard problem, and even the FTQC quantum algorithms for such tasks are merely heuristics that may fail in practice. Motivated by these challenges recent studies have focused on efficient preparation of better initial states for such tasks Berry et al. (2024).
@@
Supply chain. The final challenge to address is the cost of a quantum computer, which could be reduced by leveraging the existing semiconductor supply chain. In the final Section VIII.1, we will discuss how some of the leading fabrication, chip manufacturing, and system integrator companies, e.g., Applied Materials, Synopsys, Nvidia, and HPE, could establish a supply chain to drive down costs.
I.3 A full-stack hardware–software architecture for high-performance quantum computation

### v2

@@ I.2.3 Challenges at 10k–100k physical qubits
At the very large scale, circuit-level scaling challenges become significant (Markov, 2014, Table 1), including verification, testing, and debugging. For conventional integrated circuits, the challenge of “dark silicon” arises, where a significant fraction of the chip performs various service roles Taylor (2013). In quantum computing, FTQC creates a similar overhead.
@@
At the scale of tens of thousands of physical qubits, many fault-tolerant protocols including full-fledged magic state distillation units can be implemented and validated. Yet, the high space and time overhead of FTQC mean this scale will still fall short of demonstrating quantum utility. This prompts the need for advancements in QEC and FTQC schemes that reduce the overhead of fault tolerance, a goal actively pursued in current research trends for building “good” QECCs, that is, those with high encoding rates, such as the quantum LDPC codes Breuckmann and Eberhardt (2021).
A promising approach to reducing the FTQC overhead at this scale is adopting partially fault-tolerant compilation of quantum algorithms, for example, the compilation scheme based on error-corrected Clifford gates and space–time-efficient analog rotations Akahoshi et al. (2024); Toshio et al. (2024). This approach entirely omits the costly magic state distillation procedures that are typically required for a reliable implementation of non-Clifford gates; it also avoids decompositions of arbitrary-angle rotation gates into the Clifford$+T$ gate set via the Solovay–Kitaev algorithm or alternative techniques (see Section IV and Section A.2), which typically result in a large number of $T$ gates. Instead, arbitrary-angle rotation gates are directly executed through analog rotation and ancilla state injection followed by error detection and post-selection in a repeat-until-success fashion.
@@
Moreover, the successful realization of QEC requires low-latency integration of QPUs with GPUs, which are effective at executing a large number of identical, relatively shallow computations in parallel on different input data. Such advantages are relevant to real-time decoding, as discussed Section III.6. Another approach to relaxing the requirements of decoders is construction of better codes with faster decoding algorithms. It has been speculated that there may exist QECCs whose decoding time is independent of the code distance Gottesman (2022).
@@
Verification, testing, and debugging. As QEC circuits become more sophisticated and undergo optimization to reduce overhead, the possibilities for introducing design errors during these optimizations increase. Verification aims to catch design bugs as soon as they are introduced Viamontes et al. (2007), testing looks for problematic behaviors in a physical quantum computer (by running specific circuits) Patel et al. (2004); Maksymov et al. (2022), and debugging attempts to diagnose and correct problems nei . Historically, each of these steps became a bottleneck to scaling of classical semiconductor circuits, and required the development of new algorithmic technologies and hardware solutions to sustain scaling Lavagno et al. (2016). These tasks are much more complicated for quantum circuits than classical ones. For example, a quantum counterpart to the conventional equivalence-checking technique Yamashita and Markov (2010) must tackle unitary operators acting on exponentially large Hilbert spaces. Similarly, testing must be heavily optimized to handle the large number of trials required in view of the non-deterministic nature of quantum measurements and the frequent need of QPUs for recalibration Maksymov et al. (2020). More-sophisticated approximate testing Romanik (1997) techniques must also be developed to take the error tolerance of quantum computation into account. Finally, it is much harder to diagnose and eliminate errors Metwalli and Meter (2023); therefore, more-scalable debugging techniques are needed which can benefit from HPC hardware support.
@@ I.2.4 Challenges at 100k–1M physical qubits and beyond
Computation at extreme scales faces system-level and complexity-theoretical challenges as well as those related to physical embedding and distributed computation as quantum interconnects become a bottleneck (Markov, 2014, Table 1). For quantum computers dominated by QEC, these challenges take on specific forms, as described below. Additionally, finding killer apps for quantum computers and validating their performance remains challenging.
@@
Distributed FTQC. To achieve utility scale, tens to thousands of logical qubits are needed (Table 4). Even at a 10-4 two-qubit error rate, this translates to 1 million (or more) physical qubits, which is about an order of magnitude more than what can practically be placed in today’s dilution refrigerators (DR). Significantly larger DRs will be costly. Therefore, performing large-scale FTQC will likely require quantum interconnects between multiple DRs, which brings its own set of challenges Caleffi et al. (2024). Optical interconnects have been proposed for providing such quantum links between distinct DRs Lee et al. (2020a); Youssefi et al. (2021); Lecocq et al. (2021); Awschalom et al. (2021); Delaney et al. (2022). We discuss the first analysis of the effects of noisy optical interconnects for QECCs on superconducting computers in Section III.7 and discuss the compilation of FTQC algorithms on a multi-DR architecture in Section III.7.
Recent proposals for scalable networking of neutral-atom fault-tolerant architectures Sunami et al. (2024); Sinclair et al. (2024) suggest promising prospects for distributed quantum computing. However, while the creation of atom–photon entanglement and heralded atom–atom entanglement generation based on atom–photon interface have been demonstrated in numerous experiments, deploying entanglement among remote superconducting qubits through an interface with microwave photons is a less mature but promising direction for distributed superconducting quantum computing Storz et al. (2023).
@@
Since quantum networking technologies are still in their infancy, in Section VI, we show how adaptive circuit knitting strategies could delay the need for optical interconnects via high-performance distributed quantum evaluation of sub-circuits and classical post-processing to merge the solutions. With large-scale integration of a quantum accelerator with a classical supercomputer, dynamically dispatching workloads to and from the quantum computer will become a complex challenge. We will discuss the use of near real-time circuit synthesis and dispatching in Section V.4 on high-performance quantum–classical workload scheduling.
@@
Microarchitecture standardization. Efficient compilation and execution of large FTQC programs demands optimized and modular instruction-set architectures that are drastically different from that of conventional computers. A challenge at this scale is how to converge to an effective and versatile microarchitecture. Such microarchitectures will face additional challenges of dependability; they must be designed to be reproducible, resilient, and secure Giusto et al. (2024). Recent studies Litinski (2019); Beverland et al. (2022); Silva et al. (2024a, b) suggest that efficient instruction pipelines for FTQC using solid-state qubits comprise (potentially multi-level) magic state factories for producing high-quality resource states for consumption in a memory unit (which we call the core processor; see Section III.1). Therefore, unlike the conventional von Neumann architecture wherein data is taken from memory blocks to the gates, in the FTQC pipelines, high-quality gates are brought to the memory block. Thus, FTQC micro-architectures may better resemble that of in-memory computing technologies Verma et al. (2019). This suggests a rethinking of conventional memory hierarchies (e.g., cache, L1, and L2) for quantum memory blocks. Notably, quantum random access memory (QRAM) Giovannetti et al. (2008) is used in quantum computing literature as means for accessing and manipulating classical or quantum data in superposition. However, monumental resources are required to insure its fault tolerance Matteo et al. (2020). Quantum LDPC codes may provide a path forward for realization of feasible quantum memory blocks.
High-performance quantum operating system. Given the lengthy execution time of utility-scale FTQC algorithms and the need for frequent recalibration of QPUs, a quantum operating system (OS) must manage an excess supply of on-boarded and off-boarded DRs mid-runtime and dispatch quantum characterization, verification, and validation (QCVV) protocols on them with an appropriate cadence. This reveals a fundamental difference between FTQC compilers and the conventional compilers for classical computers: classical compilers do not require knowledge of the noise profile of the hardware, whereas the choice of QEC codes, their sizes, and ancillary modules (e.g., magic state factories) all require detailed information about the hardware noise characteristics. We describe preliminary steps toward addressing these challenges for FTQC compilers in Section III.1.
An OS for a fault-tolerant computer must therefore perform offline and real-time QPU and decoder characterization, modelling, and performance analysis, and incorporate this information into the compilation pipeline for FTQC execution. We consider three main modules for such a software suite, namely, the FTQC compiler, the emulator (including a noise profiler), and the assembler, details of which are discussed in the following section and summarized in Figure 2. An example of such a software suite is 1QBit’s TopQAD (Topological Quantum Architecture Design) toolkit 1QB Information Technologies (2024) (1QBit); Silva et al. (2024b, a). Aside from playing the role of a real-time OS for the envisioned quantum supercomputer, TopQAD is used in this paper to conduct two types of benchmarking: (a) to derive QPU and decoder performance targets, and QPU fabrication robustness targets in Section III; and (b) to perform detailed physical resource estimations for utility-scale quantum algorithms in Section IV. The compute-heavy real-time processes of the quantum OS suggest that it will ultimately be hosted on a well-integrated HPC platform communicating with the decoder and controllers in real time and refactoring the compilation/assembly of the quantum program given real-time QPU noise profiling data as depicted in Figure 2.
@@
Quantum algorithm discovery. The absence of efficient quantum memory blocks and methods for reading from and writing into them is one of the reasons for the lack of useful quantum algorithms for conventional enterprise problems involving classical big data, for which classical AI has provided major breakthroughs at an increasing pace. This read–write bottleneck eliminates the exponential speedups promised by many algorithms, for example, quantum machine learning (QML) applied to classical data. Such algorithms include quantum linear system solvers Harrow et al. (2009); Childs et al. (2017), quantum clustering Horn and Gottlieb (2001), quantum principal component analysis Lloyd et al. (2014), and quantum support vector machines Rebentrost et al. (2014). However, the promise for exponential quantum advantage still holds for quantum data Huang et al. (2022) and substantial progress has been made toward overcoming the known trainability limits of quantum neural networks Larocca et al. (2024).
@@
Moreover, processing classical data, for example, via coherent arithmetic operations, is costly for quantum computers Häner et al. (2018a, b). Indeed, the qubitized electronic-structure quantum simulation for which we provide resource estimates in Section IV uses look-up tables (coined as quantum read-only memory, or QROM) to avoid calculating trigonometric functions Babbush et al. (2018). In general, applications with classical inputs and outputs are fundamentally limited because quantum computers do not provide a universal advantage. To wit, there is provably no quantum advantage for standard comparison-based sorting—no quantum algorithm can solve the task asymptotically faster than conventional algorithms do Høyer et al. (2001).
@@
Even if a given subroutine in an important application is accelerated by a quantum oracle, it must be a single distinctive bottleneck, otherwise the impact of quantum speedup will be capped by Amdahl’s law Rodgers (1985). Shor’s algorithm for integer factorization and its variants offer a rare combination of likely exponential quantum speedup with a practical need for running the algorithm on many different inputs. In general, quantum optimization could offer a quadratic speedup Szegedy (2004) with many implicit assumptions such as lack of explicit structures in the problem instances. However, the opportunity for such quadratic speedups in practice might be slim and most likely eventually washed away at scale by the huge overhead of QEC Babbush et al. (2021a). Engineering at scale requires developing novel quantum heuristic algorithms that work in concert with their classical counterparts, see Section VII for a discussion on the possibility of accelerating classical probabilistic sampling by quantum-inspired subroutines or quantum fluctuations. Recently, there has been tremendous interest in the family of quantum approximate optimization algorithms (QAOA) Farhi et al. (2014); Blekos et al. (2024), including numerical or theoretical studies that claim potential scaling advantages might be possible Shaydulin et al. (2024); Boulebnane and Montanaro (2024); Augustino et al. (2024); Montanaro and Zhou (2024). However, small-size effects can be misleading, and contrived nature of some benchmarking tasks misrepresent scaling advantage in practical scenarios for which highly-tuned classical heuristics are available. Historically, it has proven difficult to develop quantum heuristics that remain relevant at large scales. The search for new quantum algorithms on classical inputs remains a major research avenue.
@@
Validation of quantum algorithms. There is little doubt that quantum computation can provide revolutionary advantage for quantum-mechanical problems Lloyd et al. (2014); Huang et al. (2022). But validating such algorithms—ensuring that they produce correct outputs, especially in practice—remains a challenge. For example, quantum computation of electronic spectra of molecules relies on preparation of input states with significant overlap with the ground state (whose energy is to be estimated). However, is it unclear whether the commonly adopted choices (e.g., the Hartree–Fock state) will be sufficient. Indeed, ground state preparation is a QMA-hard problem, and even the FTQC quantum algorithms for such tasks are merely heuristics that may fail in practice. Motivated by these challenges, recent studies have focused on efficient preparation of better initial states for such tasks Berry et al. (2024).
@@
Supply chain. The final challenge to address is the cost of a quantum computer, which could be reduced by leveraging the existing semiconductor supply chain. In the final Section VIII.2, we will discuss how some of the leading fabrication, chip manufacturing, and system integrator companies, for example, Applied Materials, Synopsys, Nvidia, and HPE, could establish a supply chain to reduce costs.
Overall, a natural question for the arrival of utility-scale quantum computing is the estimated timeline for development. This will be addressed next.
I.3 Exponential scaling progress: A mirage or reality?
Figure 1: Schematic plot of the number of qubits from three experiments at UCSB and Google over time, which describes a Moore’s law growth in the number of qubits. After the quantum supremacy experiment in 2019 Arute et al. (2019), a target of one million qubits at the end of the decade was projected by industry leaders. However, the current pace of hardware progress Acharya et al. (2024) suggests that goal might be postponed by several decades, assuming an optimistic scenario that none of the scaling challenges mentioned here slows or halts the progress. To arrive at utility-scale quantum computers in the 2030–2035 time frame, a major increase in the rate of progress over the next five years is needed. Our thesis is that new fabrication and systems design as well as full-stack HPC integration are required to tackle this challenge. For estimates of future scaling, we suggest using the number of qubits that can be entangled in practice (which assumes sufficient qubit connectivity with multi-qubit gates that are fast and accurate enough) Cao et al. (2023).
Since the quantum supremacy milestone Arute et al. (2019), there has been an expectation that scaling the number of qubits will follow a Moore’s law (exponential) growth over time from $\sim$50 qubits to a million qubits at the end of this decade. This goal would mark the arrival of a practical fault-tolerant quantum computer. We chart this progress in Figure 1, starting from 2014 when a repetition code experiment was performed at UCSB (nine qubits) Kelly et al. (2015), through the Google quantum supremacy on random circuits Arute et al. (2019) (53 qubits), on to the most recent surface-code error-correction experiment (105 qubits) Acharya et al. (2024). Care must be taken in plotting only the qubit quantity since this ignores other important qubit metrics such as quality, speed, and connectivity. Here, these data points represent qubit systems with some degree of consistency in these four key metrics; thus, it is a fairly reasonable timeline that has been plotted.
The trajectory of these three data points appears to follow an exponential improvement, but at a much lower slope than is needed to reach one million qubits by 2030. Larger numbers of qubits have been reported for superconducting quantum processors, but even for these more optimistic characterizations of functional qubits, the scaling falls short of the original industry expectation. Will it be possible to greatly increase the slope over the next five years, corresponding to a double exponential growth? We think this is unlikely because the added challenges for scaling beyond 1000 qubits—as detailed here—could hinder even staying on the current growth path.